(a) Solve the initial value problem 2 +2r + r = 8(t - 2), z(0)=1, 20) = 2 (b) Consider the initial value problem -2 -5 z(0
Answer #1

Aven (4-2)$ = = a + اه را= (6) X (= (ه) * on both sides + (10w408 مقهاها Apply (2-t) } ا =143 + 213 + { }L و (xb 3 - 433اوه +u(t-2) و2 م كات ( (t-2) 8 (4-2) NOW IS ı (3+1)? i{ csovs ] { tét \{ stij}a te t{ is ve} = (174) }- cop tët) - - ēt (t-1) .